OpenStudy (anonymous):

I have to solve the system of equations in this problem (which i have noooo idea what that means XD) 2x + 3y = 1 y = 3x = 15 help??

jimthompson5910 (jim_thompson5910):

there seems to be a typo in "y = 3x = 15"

OpenStudy (anonymous):

Oh! Its supposed to be y = 3x + 15 oops XD

jimthompson5910 (jim_thompson5910):

start with the first equation 2x + 3y = 1 and replace every y with 3x+15 (since y = 3x+15)

jimthompson5910 (jim_thompson5910):

doing that gives you 2x + 3y = 1 2x + 3(3x+15) = 1 solve for x from here

OpenStudy (anonymous):

oh!

OpenStudy (anonymous):

Wow thanks @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you're welcome

jimthompson5910 (jim_thompson5910):

and once you get the value of x, you use y = 3x+15 to find y

OpenStudy (anonymous):

ok cool

OpenStudy (anonymous):

so @jim_thompson5910 the correct answer for x is -1?

jimthompson5910 (jim_thompson5910):

x = -1 is incorrect

OpenStudy (anonymous):

Oh really?

OpenStudy (anonymous):

I did 2x + 3(3x + 15) = 1 2x + 9x + 45 = 1 11x + 45 = 1 11x = -44 x = -1 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

11x = -44 does not lead to x = -1

OpenStudy (anonymous):

OOOH! its -4 ? @jim_thompson5910

OpenStudy (anonymous):

oops XD

jimthompson5910 (jim_thompson5910):

yep, x = -4

jimthompson5910 (jim_thompson5910):

and y = ??

OpenStudy (anonymous):

y = 2?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

y = 3x + 15 y = 3(-4) + 15 .... replace x with -4 y = -12 + 15 y = 3

jimthompson5910 (jim_thompson5910):

you were close

OpenStudy (anonymous):

oh i just added wrong XD thsnkyou!!

jimthompson5910 (jim_thompson5910):

you're welcome